Circular No. 5785

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GRS 0834–43 S. Brandt, A. J. Castro-Tirado, and N. Lund, Danish Space Research

Institute, report: “The x-ray transient GRS 0834–43 (GS 0834–430; IAUC 5122, 5139, 5142, 5145, 5180) has been observed by the WATCH all-sky monitor on EURECA during May 4.3-5.1 and 5.5-7.1 UT. The flux from the source averaged 200 ± 50 mCrab at 6-20 keV. Pulsations were not detected above 50 mCrab in the pulsed fraction. During observation periods Feb. 3-13, Mar. 14-Apr. 10, and Apr. 25-26, the source was not detected with a limit of 50 mCrab. The 0^{o}.5-radius error circle (3σ) of the WATCH detection does not rule out MX 0836–42 as the source of the x-ray flux.”

SUPERNOVA 1993J IN NGC 3031 J. M. Marcaide, Universitat de Valencia; A. E. E. Rogers, Haystack

Observatory; A. Alberdi, J. C. Guirado, and A. Rius, Consejo Superior de Investigaciones Científicas, Spain; P. Elósegui and I. I. Shapiro, Harvard-Smithsonian Center for Astrophysics; E. Pérez, Instituto de Astrofísica de Canarias; and A. R. Whitney, Haystack Observatory, report: “Using an interferometer composed of NASA’s Deep Space Network 70-m antenna in Madrid, Spain, and the Haystack Observatory’s 18-m antenna in Westford, MA, at a wavelength of 3.6 cm during Apr. 24.90-25.06 UT, we detected SN 1993J. Fringe spacing was ∼ 0′′.0014. We used Mark III recording equipment, synthesized a 64-MHz band, and correlated the data at the Haystack Observatory. This first VLBI detection of SN 1993J yields a size (FWHM for a circular gaussian source model) of 0′′.00025 ± 0′′.0001, assuming a total flux density of 13.4 mJy (IAUC 5775).”

Corrigendum. On IAUC 5780, line 13, for (see also IAUC 5577).” read (see also IAUC 5777).”

V1974 CYGNI G. F. Lawrence, R. D. Gehrz, and K. J. Draeger report infrared ob-

servations of V1974 Cyg (N Cyg 1992) obtained at the University of Minnesota’s O’Brien Observatory 0.76-m telescope (+ InSb detector): Apr. 21.3 UT, H = 11.55 ± 0.08, K = 10.62 ± 0.05, L′ [3.8 μm] = 7.10 ± 0.15. The beam size was 27′′ and the throw between the source and reference beams was 40′′.

1993 May 8 Daniel W. E. Green